OpenStudy (rsadhvika):

When the automobile manufacturer releases a new model in its WOW series, it promises that drivers can expect to get mileage that varies by no more than 6.5 miles per gallon from its field-tested average of 28.4 miles per gallon. Which solution represents the range of mileage you can expect from this car?

OpenStudy (rsadhvika):

I know the answer is \(28.4 \pm 6.5\)

OpenStudy (rsadhvika):

But I am stuck at solving this problem by using an absolute value equation... can we turn the given info to an absolute equaiton and sovle ?

OpenStudy (rsadhvika):

like \(28.4 = |x| + ... \) something

OpenStudy (rsadhvika):

I feel somehow i forgot the basics :(

OpenStudy (whpalmer4):

|28.4-mileage| <= 6.5 test it out: if the mileage achieved is the lower bound (28.4-6.5 = 21.9), |28.4-21.9| = | 6.5| = 6.5 <= 6.5 if the mileage achieved is the upper bound (28.4+6.5 = 34.9), |28.4-34.9| = | -6.5| = 6.5 <= 6.5

OpenStudy (whpalmer4):

You know that the magnitude of the predicted mileage - actual mileage is the variation, right? So our equation is the same as saying 28.4-6.5 <= mileage <= 28.4+6.5 and you should be able to convince yourself that the two forms are equivalent...

OpenStudy (rsadhvika):

thanks a lot ! i get it completely now xD